OpenStudy (anonymous):

what is the vertex form for:y=-2x^2+12x-18 ?

OpenStudy (anonymous):

y = -2(x^2 - 6x) - 18 y = -2(x^2 - 6x + 9) - 18 + 18 y = -2(x - 3)^2 vertex at (3, 0) is a maximum with y-axis as the axis of symmetry.

OpenStudy (anonymous):

The "+ 0" is supposed to be left out and is "understood".

OpenStudy (anonymous):

Sorry, x = 3 is axis of symmetry. All else holds.

OpenStudy (anonymous):

Making sense to you now?